We are interested in n = 6 devices. Note that each device constitutes a trial with one of two outcomes: device functions at least 15 hours or it doesn't. Each of the 6 trials are Bernoulli trials and we assume the trials are independent. We also have a constant probability of success p = 2 3 . We use the binomial model.
Answer #1

solution P= 2/5 n=6 P673) - P(x-3) + P(x=4) + P(x=5) + P(x=6) = 63 0.6667)3 (0:333)34 66 66667)*60.333)? +605(0-66667)5(0:333